A data set about speed dating includes like ratings of male dates made by the female dates. The summary statistics are n= 195, x= 5.88, s= 2.06. Use a 0.05 significance level to test the claim that the population mean of such ratings is less than 6.00. Assume that a simple random sample has been selected. Identify the null and alternative hypotheses, test statistic, P.value, and state the final conclusion that addresses the original claim. What are the null and alternative hypotheses? O A. Ho-p=6.00 OB. Ho-= 6.00 H - p+6.00 Hy: <6.00 OC. H, <6.00 OD. Ho: p= 6.00 H, p>6.00 Hy: p> 6.00 Determine the test statistic. (Round to two decimal places as needed.). Determine the P-value. (Round to three decimal places as needed.) State the final conclusion that addresses the original claim. Ho There is V evidence to conclude that the mean of the population of ratings is TA 76.00. alert your answerist
Answer #1

Here,

Sample mean \(=\bar{x}=5.88\)

Sample \(S . D=s=2.06\)

Sample size \(=n=195\)

Degrees of freedom \(=194\)

Significance Level \(=\alpha=0.05\)

Step 1: \(H_{0}: \mu=6\)

\(H_{1}: \mu<6\) (Left tailed test)

Step 2: The test statistic is, \(t=\frac{\bar{x}-\mu}{\left(\frac{s}{\sqrt{n}}\right)}=-0.81\)

Step \(3: P-\) value \(=P(t<-0.81)=0.209\)

Step \(4:\) We fail to reject \(H_{0} .\) There is not sufficient evidence to conclude that the mean of the population of ratings is less than \(6.00\).